OpenStudy (anonymous):

Power series problem: n!(2x-1)^n from 1 to infinity

OpenStudy (abb0t):

are you being asked to determine whether the power series converges or not?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when simplified i got absolute value of (2x-1) Lim n--->00 n+1

OpenStudy (abb0t):

use ratio test. works about 75-80% of the time for most power series. and best test to use when you have factorials.

OpenStudy (anonymous):

i did that, but not sure if i am right..

OpenStudy (abb0t):

\[L = \lim_{n \rightarrow \infty}\left| \frac{ a_{n+1} }{ a_n } \right|\] if L <1 absolutely convergent, and thus convergent L > 1 is divergent L = 1 use different test. but it SHOULD work.

OpenStudy (abb0t):

\[L = \lim_{n \rightarrow \infty}\left| \frac{ (n+1)! }{ (2x-1)^{n+1} } \times \frac{(2x-1)^n}{n!} \right| = \lim_{n \rightarrow \infty }\left| \frac{ n(n+1)}{ (2x-1)(2x-1)^n } \ \times \frac{ (2x-1)^n }{ n } \right|\]

OpenStudy (abb0t):

I'm sure you can finish it from here.

OpenStudy (anonymous):

so when simplified should be (n+1)(2x-1) right?

OpenStudy (anonymous):

so i pull the 2x-1 out and im left with the Lim n--->00 of n+1 ?

OpenStudy (abb0t):

yep.

OpenStudy (anonymous):

thanks abbot!!

OpenStudy (anonymous):

Your series diverges everywhere except for x =1/2. You can do it by the divergence test, the nth terms does not go to zero, then the series diverges

OpenStudy (anonymous):

hey Elias, how did you figure that out?

OpenStudy (anonymous):

It is a power series. if \( x\ne \frac 1 2\), then n!(2x-1)^n goes to infinity with n, so the nth-term does not go to zero, so the series is divergent.